Please, please help? I need to solve:

by addition: 5x - 3y = 13 4x - 3y = 11
by graphing:x - y = 5
x + y = -1






















Answer by checkley71(8403)

5X-3Y=13
4X-3Y=11 NOW SUBTRACT
--------------------
5X-4X=13-11
X=2 ANSWER.
5*2-3Y=13
10-3Y=13
-3Y=13-10
-3Y=3
Y=3/-3
Y=-1 ANSWER.
PROOF.
4*2-3*-1=11
8+3=11
11=11